### v1

# PDLP: A Practical First-Order Method for Large-Scale Linear Programming

Abstract
We present PDLP, a practical first-order method for linear programming (LP) designed to solve large-scale LP problems. PDLP is based on the primal-dual hybrid gradient (PDHG) method applied to the minimax formulation of LP. PDLP incorporates several enhancements to PDHG, including diagonal preconditioning, presolving, adaptive step sizes, adaptive restarting, and feasibility polishing. Our algorithm is implemented in C++, available in Google’s open-source OR-Tools library, and supports multithreading.
To evaluate our method, we introduce a new collection of eleven large-scale LP problems with sizes ranging from 125 million to 6.3 billion nonzeros. PDLP solves eight of these instances to optimality gaps of 1% (with primal and dual feasibility errors of less than $10^{-8}$) within six days on a single machine. We also compare PDLP with Gurobi barrier, primal simplex, and dual simplex implementations. Gurobi barrier solves only three instances, exceeding our 1TB RAM limit on the other eight. While primal and dual simplex are more memory-efficient than the barrier method, they are slower and solve only three instances within six days.
Compared with the conference version of this work (in: Advances in Neural Information Processing Systems 34 (NeurIPS 2021)), the key new contributions are: (i) feasibility polishing, a technique that quickly finds solutions that are approximately optimal but almost exactly feasible (without which only three of the eleven problems can be solved); (ii) a multithreaded C++ implementation available in Google OR-Tools; and (iii) a new collection of large-scale LP problems. Note that the conference version should be referred to for comparisons with SCS and ablation studies, which we do not repeat in this paper.
1 Introduction
Applications of linear programming (LP) are widespread and span many different fields [17]. Simplex [61] and interior-point methods [86] are the standard approaches to solving LP problems. They are widely used in practice and are known for their reliability and ability to solve instances to high numerical precision. However, these methods face scaling limitations when applied to extremely large problems. For interior point methods, the limiting factor is often the memory required by sparse matrix factorizations. While the simplex method is less likely to run out of memory because it factorizes a smaller matrix, it often requires prohibitively many iterations. Moreover, it is difficult to adapt these methods to exploit highly-parallel modern computing hardware such as multithreading, distributed computing, and graphics processing units (GPUs).
A parallel stream of research that has had relatively little impact on the practice of LP, first-order methods (methods that use function and gradient information) have undergone an intense period of study over the past twenty years [32, 13, 62, 24, 71, 45], resulting in incredible success in machine learning [52, 50]. The benefit of these methods is that they have cheap iterations that can readily utilize highly parallel modern computing hardware and quickly obtain approximately optimal solutions, which are sufficient for most large-scale machine learning applications. Building on this literature, this paper enhances an existing first-order method, primal-dual hybrid gradient (PDHG), popularized by Chambolle and Pock [24] in the computer vision community, to develop a specialized first-order method for LP.
Despite their appealing computational properties, a longstanding barrier to adopting first-order methods for LP is their inability to reliably solve instances to high numerical precision in reasonable timeframes. For LP, it’s especially critical to find feasible solutions, both primal and dual. Even small violations of primal constraints (for instance, flow conservation) are difficult for practitioners to interpret. Similarly, finding dual feasible solutions is just as essential because of LP’s use for providing bounds on the optimal objective value such as within branch-and-bound methods [85]. As a concrete example of the risks associated with an inaccurately solved LP problem, consider the ‘irish-electricity’ instance, an LP instance based on a unit commitment case study of the Irish electric grid [63, 23], which has been used in Mittelmann’s LP benchmark [64]. On this instance, SCS 3.27 [72], a state-of-the-art first-order method with its default termination tolerance of $10^{-4}$, terminates in 40 seconds. However, the objective value for the returned solution is $2.21\times 10^{6}$, which is off from the true optimal objective value [^1] Found using Gurobi with crossover enabled. of $2.56\times 10^{6}$ by more than 13%. On the other hand, SCS with a tighter termination tolerance of $10^{-8}$ does not terminate successfully within an hour. This issue—where moderate feasibility violations lead to such undesirable behavior, which is only detectable using solutions provided from a more accurate solver—justifies practitioners’ caution in adopting first-order methods for LP.
To tackle this reliability issue, we apply a series of enhancements to PDHG including: adaptive restarting, dynamic primal-dual step size selection, and diagonal preconditioning. We call this enhanced PDHG algorithm, PDLP (PDHG for LP). The impact of these enhancements was demonstrated in the conference version of this paper [7]. For example, for finding high accuracy solutions on a standard LP benchmark, NETLIB [37], our enhancements enabled a 75×75$\times$75 × speed-up over PDHG without any enhancements and 19×19$\times$19 × speed-up [7, Table 24] over SCS [72], a first-order method for conic optimization.
Compared with the conference version of this paper [7], the new contributions of this paper include:
  1. 1.
    An implementation of PDLP in multithreaded C++ code which is now part of Google’s open-source OR-Tools package (https://developers.google.com/optimization). The conference version of our paper uses single-thread Julia code [^2] The prototype Julia code, FirstOrderLp.jl, is available at https://github.com/google-research/FirstOrderLp.jl..
  2. 2.
    A new collection of large-scale LP instances with between 125 million and 6.3 billion nonzeros to benchmark our method and a set of experiments demonstrating favorable results compared to Gurobi, a state of the art commerical solver. In comparison, the conference paper focused on smaller standard benchmark instances: the largest problem tested had 90 million nonzeros.
  3. 3.
    An additional enhancement that allows our algorithm to rapidly find solutions that are feasible but only approximately optimal. Our computational results demonstrate this new heuristic is extremely effective for these large-scale problems.
Note that this paper does not subsume the conference version because we do not repeat its numerical experiments. These experiements were:
  1. 1.
    An ablation study that individually disabled each enchancement and also compared with reasonable alternatives.
  2. 2.
    Starting from a naive PDHG implementation, a demonstration of the incremental impact of each enchancement.
  3. 3.
    Comparisions with existing solvers (SCS and Gurobi) on instances from medium-size LP benchmarks and a large-scale LP example derived from the PageRank problem [19].
Given the C++ code is the exclusive focus of this paper, we decided neither to repeat results verbatim from the conference paper nor rerun these experiments using the C++ code as the conclusions would largely be the same. For reference [56, Table 8] shows that the single threaded C++ code achieves roughly a 1.5×1.5$\times$1.5 × speed-up over the Julia code.
Outline
Section 2 introduces the PDHG algorithm and the notation we use throughout the paper. Section 3 explains the practical improvements we have added to the PDHG algorithm. Section 4 introduces our new feasibility polishing algorithm. Section 5 describes the C++ implementation. Section 6 describes the collection of large-scale instances. Section 7 presents our numerical results.
1.1 Related literature
Classic methods for linear programming.
The simplex method was the first successful algorithm for LP, having been developed by Dantzig and others in the late 1940s, with improvements over the subsequent decades as the economic importance of LP grew [61]. The simplex method begins at a vertex and pivots between vertices, monotonically improving the objective function with each step. Interior-point methods [86] were a major breakthrough in the 1980s, coupling much stronger theoretical guarantees with practically successful implementations. In contrast to the simplex method, interior-point methods (IPMs) traverse the interior of the feasible region, progressing toward optimality by following a so-called central path [77]. Commercial and open-source solvers such as Gurobi, CPLEX, COPT, and HiGHS leverage both of these methods, often in combination, to provide reliable and accurate solutions to LP problems.
PDHG.
The primal-dual hybrid gradient (PDHG) method, originally proposed in [91, 35], has been extended and analyzed by numerous researchers [35, 76, 24, 44, 27, 26, 5]. PDHG is a form of operator splitting [12, 81], and it is closely related to the alternating direction method of multipliers (ADMM) and Douglas-Rachford splitting (DRS) [34, 18, 70], all of which can be interpreted as preconditioned instances of the proximal point method (PPM) [79, 34, 73, 57]. However, unlike ADMM, DRS, or PPM, when applied to LP problems, PDHG is matrix-free, relying solely on matrix-vector multiplications.
PDLP.
The conference version of this work [7] introduced PDLP, a restarted PDHG with several practical enhancements specialized for solving LP problems. Experiments were based on a prorotype Julia implementation. This paper includes an additional enhancement, and the experiments discussed in this paper are based on a C++ implementation of PDLP, which has been open-sourced through Google OR-Tools since 2022 [74]. More recently, a GPU implementation, dubbed cuPDLP, demonstrated strong numerical performance comparable to commercial LP solvers, garnering attention from both the academic community and the solver industry [56, 60]. Variants of (cu)PDLP have been incorporated into both commercial and open-source optimization solvers, including COPT [38], FICO Xpress [88], HiGHS [49], and NVIDIA cuOpt [16].
The success of PDLP has spurred theoretical analysis. Lu and Yang [55] proved that PDHG exhibits a linear convergence rate when applied to LP. Applegate et al. [8] showed that a faster linear convergence rate can be achieved through restarts, with the rate matching the worst-case lower bound. Applegate et al. [9] demonstrated that for infeasible or unbounded LP instances, PDHG iterates diverge toward a certificate of infeasibility or unboundedness, allowing for infeasibility detection without additional computation. Lu and Yang [58] established that PDHG displays a two-stage behavior when applied to LP: in the first stage, it identifies the optimal active variables in only finitely many iterations; in the second stage, it solves a homogeneous linear inequality system at an improved linear rate. This two-stage behavior helps explain the slow convergence observed when an LP instance is near-degenerate. Hinder [47] shows an iteration complexity bound for restarted PDHG applied to totally unimodular LP of $\tilde{O}(Hm^{2.5}\sqrt{\text{nnz}(A)})$ where $H$ is the largest absolute value of the right-hand side and objective vector, $m$ is the number of rows and $\text{nnz}(A)$ the number of nonzeros in the constraint matrix. Unlike other results, this bound does not involve any condition numbers. Zikai and Freund [87] introduced the concept of level-set geometry, showing that PDHG converges more quickly for LP instances with regular, non-flat level sets around the optimal solution. Recently, additional numerical improvements have been proposed to further accelerate the algorithm. In particular, Zikai and Freund [87] introduced a central-path-based preconditioner for PDHG that significantly improves convergence. Further, Lu and Yang [59] incorporated the Halpern iteration [43] into PDLP, achieving accelerated theoretical guarantees and improved practical performance.
Other first-order methods for LP
Lan, Lu, and Monteiro [53] and Renegar [78] developed first-order methods (FOMs) for LP as a particular case of semidefinite programming with sublinear convergence rates. The FOM-based solvers apply to more general problem classes like cone or quadratic programming. In contrast, some of the enhancements that constitute PDLP are specialized, either in theory or practice, for LP (namely restarts [8] and presolving). A number of authors [84, 33, 39, 89, 65] have proposed linearly convergent FOMs for LP; however, to our knowledge, none have been the subject of a comprehensive computational study. FOM-based software packages that can be used to solve LP include the following:
  • SCS [68, 67] is an ADMM-based solver for convex quadratic cone programs implemented in C. SCS applies ADMM [18] to the homogeneous self-dual embedding of the problem [90], which yields infeasibility certificates when appropriate. One of the key computational challenges of ADMM-based methods is solving a linear system at every iteration. SCS provides two options to tackle this subroutine: a direct solver, which computes and reuses the factorization of the matrix, and an indirect solver, which uses conjugate gradient methods to iteratively solve these linear systems, offering scalability trade-offs.
  • OSQP [82] is another ADMM-based solver for convex quadratic programming implemented in C. Compared to SCS, a major difference is that OSQP applies ADMM to the original problem rather than the homogeneous self-dual embedding. It also has direct and indirect options for solving the linear systems at each iteration. A related solver, COSMO [36], uses the same algorithm for convex quadratic cone programs and is implemented in Julia.
  • ECLIPSE [11] is a distributed LP solver that employs accelerated gradient descent to solve a smoothed dual formulation of LPs. It is specifically designed to handle large-scale LPs with particular decomposition structures, such as those encountered in web applications.
  • ABIP and ABIP+ [54] are new interior-point method (IPM) solvers for cone programming. ABIP uses a few steps of ADMM to approximately solve the sub-problems that arise when applying a path-following barrier algorithm to the homogeneous self-dual embedding of the problem. Although IPMs are technically second-order methods, the use of ADMM in the inner loop allows the solver to scale to larger problems than traditional IPMs can handle. ABIP+ [30] adapts many of the improvements from PDLP to ABIP to improve its performance.
  • TFOCS [14] is a MATLAB solver designed for solving a variety of convex cone problems that arise in areas such as signal processing, machine learning, and statistics, including applications like the Dantzig selector, basis pursuit denoising, and image recovery. In principle, TFOCS can also be applied to LP. It utilizes a variant of Nesterov’s accelerated method [66] to solve a smoothed version of the primal-dual conic problem.
  • ProxQP [10] is a solver specifically designed for convex quadratic programming (QP), particularly instances arising in real-time robotics applications. It uses a primal-dual proximal method of multipliers and requires solving a linear system at each iteration through Cholesky factorization. ProxQP can also leverage warm-start strategies, which is especially beneficial for robotics applications.
2 Preliminaries
In this section, we introduce the notation we use throughout the paper, summarize the LP formulations we solve, and introduce the baseline PDHG algorithm.
Notation.
Let ${\mathbb{R}}$ denote the set of real numbers, ${\mathbb{R}}^{+}$ the set of nonnegative real numbers, and ${\mathbb{R}}^{-}$ the set of nonpositive real numbers. Let ${\mathbb{N}}$ denote the set of natural numbers (starting from one). We endow ${\mathbb{R}}^{n}$ with the standard dot product $\langle x,y\rangle=x^{\top}y$ and its induced norm $\|\cdot\|_{2}.$ We will also use $\|\cdot\|_{p}$ to denote the $\ell_{p}$ norm for a vector. The $\ell_{\infty}$-norm, $\|v\|_{\infty}$, computes the maximum absolute value of the vector $v.$ For a vector $v\in{\mathbb{R}}^{n}$, we use $v^{+}$ and $v^{-}$ for their positive and negative parts, i.e., $v_{i}^{+}=\max\{0,v_{i}\}$ and $v_{i}^{-}=-\min\{0,v_{i}\}$. The symbol $v_{1:m}$ denotes the vector with the first $m$ components of $v$. The symbol $\mathbf{1}$ denotes the vector of all ones. Given a matrix $A\in{\mathbb{R}}^{m\times n},$ the symbols $A_{i,\cdot}$ and $A_{\cdot,j}$ correspond to the $i$th column and $j$th row of $A$, respectively. We use $\|\cdot\|_{2}$ to denote the spectral norm for a matrix, i.e., its top singular value.
Let $\mathcal{X}\subseteq{\mathbb{R}}^{n}$ be a closed convex set. The orthogonal projection onto $\mathcal{X}$ is given by
$\mathop{\bf proj}_{\mathcal{X}}(x)=\mathop{\rm argmin}_{w\in\mathcal{X}}\|x-w\|_{2}.$
Define the mapping $p\colon{\mathbb{R}}^{n}\times({\mathbb{R}}\cup\{-\infty\})^{n}\times({\mathbb{R}}\cup\{\infty\})^{n}\to{\mathbb{R}}\cup\{\infty\}$ given by
$p(y;\ell,u):=u^{\top}y^{+}-\ell^{\top}y^{-}.$
Given two vectors ({}$\ell\in({\mathbb{R}}\cup\{-\infty\})^{n}$ and $u\in({\mathbb{R}}\cup\{\infty\}\})^{n}$ satisfying $\ell\leq u.$ The function $p(\cdot;\ell,u)$ is piecewise linear and convex. With a slight abuse of notation, we will also use $p$ to denote an analogous function taking $m$-dimensional inputs as implied by the context.
Linear Programming.
We solve primal-dual LP problems of the form [83]:
$\mathop{\rm minimize}_{x\in{\mathbb{R}}^{n}}~{}~{}&~{}c^{\top}x&\\ \text{subject to:}~{}~{}&~{}\ell_{c}\leq Ax\leq u_{c}\\ &~{}\ell_{v}\leq x\leq u_{v},\end{aligned}\qquad\qquad\qquad\begin{aligned} \mathop{\rm maximize}_{y\in{\mathbb{R}}^{m},r\in{\mathbb{R}}^{n}}\quad&-p(-y;\ell_{c},u_{c})-p(-r;\ell_{v},u_{v})\\ \text{subject to:}\quad&c-A^{\top}y=r\\ &y\in\mathcal{Y}\\ &r\in\mathcal{R}.\end{aligned}$ (1)
where $A\in{\mathbb{R}}^{m\times n}$, $c\in{\mathbb{R}}^{n}$, $\ell_{c}\in({\mathbb{R}}\cup\{-\infty\})^{m}$, $u_{c}\in({\mathbb{R}}\cup\{\infty\})^{m}$, $\ell_{v}\in({\mathbb{R}}\cup\{-\infty\})^{n}$, $u\in({\mathbb{R}}\cup\{\infty\})^{n}$, and the sets $\mathcal{Y}\subseteq{\mathbb{R}}^{m}$ and $\mathcal{R}\subseteq{\mathbb{R}}^{n}$ are Cartesian products with their $i$th components given by
$&\mathcal{Y}_{i}:=\begin{cases}\{0\}&(\ell_{c})_{i}=-\infty,~{}(u_{c})_{i}=\infty,\\ {\mathbb{R}}^{-}&(\ell_{c})_{i}=-\infty,~{}(u_{c})_{i}\in{\mathbb{R}},\\ {\mathbb{R}}^{+}&(\ell_{c})_{i}\in{\mathbb{R}},~{}(u_{c})_{i}=\infty,\\ {\mathbb{R}}&\text{otherwise};\end{cases}\quad\text{ and }\qquad\end{aligned}\begin{aligned} &\mathcal{R}_{i}:=\begin{cases}\{0\}&(\ell_{v})_{i}=-\infty,~{}(u_{v})_{i}=\infty,\\ {\mathbb{R}}^{-}&(\ell_{v})_{i}=-\infty,~{}(u_{v})_{i}\in{\mathbb{R}},\\ {\mathbb{R}}^{+}&(\ell_{v})_{i}\in{\mathbb{R}},~{}(u_{v})_{i}=\infty,\\ {\mathbb{R}}&\text{otherwise};\end{cases}\end{aligned}$ (2)
which corresponds with the set of variables $(y,r)$ for which the dual objective is finite. We highlight that the formulation we consider (1) is slightly different from the formulation in the conference version of this paper [7], which only considers a lower bound $\ell_{c}\leq Ax$ and has additional equality $Gx=q$.
This pair of primal-dual problems is equivalent to the minimax problem [83]:
$\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}\mathcal{L}(x,y):=c^{\top}x+y^{\top}Ax-p(y;-u_{c},-\ell_{c})$ (3)
with $\mathcal{X}:=\{x\in{\mathbb{R}}^{n}:\ell_{v}\leq x\leq u_{v}\}$. Notice that the variable $r$ in (1) does not appear explicitly in the minimax formulation. This absence leaves an algorithmic choice for recovering the reduced costs, $r$, which impacts when the termination criteria is satisfied but does not affect the iterates themselves. A natural choice is to set $r=\mathop{\bf proj}_{\mathcal{R}}(c-A^{\top}y)$; we discuss a more advanced choice in Appendix A.
PDHG.
When specialized to (3), the PDHG algorithm [24] takes the form:
$\begin{split}x^{k+1}&=\mathop{\bf proj}_{\mathcal{X}}(x^{k}-\tau(c-A^{\top}y^{k}))\\ y^{k+1}&=y^{k}-\sigma A\left(2x^{k+1}-x^{k}\right)-\sigma\mathop{\bf proj}_{[-u_{c},-\ell_{c}]}\left(\sigma^{-1}y^{k}-A\left(2x^{k+1}-x^{k}\right)\right)\end{split}$ (4)
where $\tau,\sigma>0$ are primal and dual step sizes, respectively. A quick case-by-case argument shows $y^{k+1}\in\mathcal{Y}$. For instance, suppose the $i$-th coordinate bounds satisfy $(\ell_{c})_{i}=-\infty$ and $(u_{c})_{i}\in\mathbb{R}$, and let $\widehat{y}^{k}\;=\;y^{k}\;-\;\sigma\,A\left(2\,x^{k+1}\;-\;x^{k}\right).$ Then, we have
$y_{i}^{k+1}\;=\;\begin{cases}0&\text{if }\widehat{y}_{i}^{k}\;\geq\;-\sigma\,(u_{c})_{i},\\ \widehat{y}_{i}^{k}\;+\;\sigma\,(u_{c})_{i}&\text{otherwise},\end{cases}$
and, thus, in any case $y_{i}^{k+1}\in{\mathbb{R}}^{-}=\mathcal{Y}_{i}$, as we wanted. The other cases follow analogously.
PDHG is known to converge to an optimal solution when $\tau\sigma\|A\|_{2}^{2}<1$ [27, 25]. We reparameterize the step sizes by
$\tau=\eta/\omega\quad\text{and}\quad\sigma=\omega\eta\qquad\text{with }\eta\in(0,\infty)\quad\text{and}\quad\omega\in(0,\infty).$ (5)
We call $\omega\in(0,\infty)$ the primal weight, and $\eta\in(0,\infty)$ the step size. Under this reparameterization, PDHG converges for all $\eta<1/\|A\|_{2}$. This parameterization allows us to control the scaling between the primal and dual iterates with a single parameter $\omega$. From now on and for the rest of the paper, we will use $z$ as a placeholder for the concatenation of primal dual vectors $(x,y).$ We use the term primal weight to describe $\omega$ because it is the weight on the primal variables in the following norm:
$\|z\|_{\omega}:=\sqrt{\omega\|x\|_{2}^{2}+\frac{\|y\|_{2}^{2}}{\omega}}.$
This norm plays a role in the theory for PDHG [25] and later algorithmic discussions.
3 Practical algorithmic improvements
In this section, we briefly describe algorithmic enhancements for PDHG that were introduced in the conference version of this paper [7]. We omit a discussion of presolve because we do not use it for the numerical experiments in this paper; see also the mention of presolve in Section 5. A separate section (Section 4) is reserved to introduce feasibility polishing, which is new to this work.
Note that while our enhancements are inspired by theory, our focus is on practical performance. The algorithm as a whole does not have convergence guarantees. Guarantees for individual enhancements are noted at the end of the section.
Algorithm 1 presents pseudo-code for PDLP. We modify the step sizes, add restarts, dynamically update the primal weights, and use multithreading. Before running Algorithm 1 we apply diagonal preconditioning.
Periodic checks of restarts and termination criteria
There are some minor differences between the pseudo-code and the actual code. In particular, we only evaluate the restart or termination criteria (Line 1) every $64$ iterations. This reduces the associated overhead of these evaluations with minimal impact on the total number of iterations. We also check the termination criteria before beginning the algorithm and if we detect a numerical error.
1Input: An initial solution $z^{0,0}$;
2 Initialize outer loop counter $n\leftarrow 0$, total iterations $k\leftarrow 0$, step size $\hat{\eta}^{0,0}\leftarrow 1/\|A\|_{\infty}$, primal weight $\omega^{0}\leftarrow$ InitializePrimalWeight(c, q);
3 repeat
4       $t\leftarrow 0$;
5       repeat
6             $z^{n,t+1},\eta^{n,t+1},\hat{\eta}^{n,t+1}\leftarrow$ AdaptiveStepOfPDHG($z^{n,t},\omega^{n},\hat{\eta}^{n,t},k$) ;
7             $\bar{z}^{n,t+1}\leftarrow\frac{1}{\sum_{i=1}^{t+1}\eta^{n,i}}\sum_{i=1}^{t+1}\eta^{n,i}z^{n,i}$ ;
8             $z_{\text{c}}^{n,t+1}\leftarrow$ GetRestartCandidate($z^{n,t+1},\bar{z}^{n,t+1},z^{n,0}$) ;
9             $t\leftarrow t+1$, $k\leftarrow k+1$ ;
10
11      until restart or termination criteria holds;
12       restart the outer loop. $z^{n+1,0}\leftarrow z_{\text{c}}^{n,t}$, $n\leftarrow n+1$;
13       $\omega^{n}\leftarrow$PrimalWeightUpdate($z^{n,0},z^{n-1,0},\omega^{n-1}$) ;
14
15until termination criteria holds;
Output: $z^{n,0}$.
Algorithm 1 PDLP (after preconditioning and presolve)
1Function AdaptiveStepOfPDHG ($z^{n,t}$, $\omega^{n}$, $\hat{\eta}^{n,t},k$):
2       $(x,y)\leftarrow z^{n,t}$, $\eta\leftarrow\hat{\eta}^{n,t}$ ;
3       for $i=1,\dots,\infty$ do
4             $x^{\prime}\leftarrow\mathop{\bf proj}_{\mathcal{X}}(x-\frac{\eta}{\omega^{n}}(c-A^{\top}y))$ ;
5             $y^{\prime}\leftarrow y-A(2x^{\prime}-x)-\eta\omega^{n}\mathop{\bf proj}_{[\ell_{c},u_{c}]}\left((\eta\omega^{n})^{-1}y-A(2x^{\prime}-x)\right)$;
6             $\bar{\eta}\leftarrow\frac{\|(x^{\prime}-x,y^{\prime}-y)\|_{\omega^{n}}^{2}}{2(y^{\prime}-y)^{\top}A(x^{\prime}-x)}$ ;
7             $\eta^{\prime}\leftarrow\min\left((1-(k+1)^{-0.3})\bar{\eta},(1+(k+1)^{-0.6})\eta\right)$ ;
8             if $\eta\leq\bar{\eta}$ then
9                   return $(x^{\prime},y^{\prime})$, $\eta$, $\eta^{\prime}$
10             end if
11            $\eta\leftarrow\eta^{\prime}$ ;
12
13       end for
14
15
Algorithm 2 One step of PDHG using our step size heuristic
Step size choice.
The convergence analysis [25, Equation (15)] of standard PDHG (4) relies on a small constant step size
$\eta\leq\frac{\|z^{k+1}-z^{k}\|_{\omega}^{2}}{2(y^{k+1}-y^{k})^{\top}A(x^{k+1}-x^{k})}$ (6)
where $z^{k}=(x^{k},y^{k})$. Classically one would ensure (6) by picking $\eta=\frac{1}{\|A\|_{2}}$. This is overly pessimistic and requires estimation of $\|A\|_{2}$. Instead our AdaptiveStepOfPDHG adjusts $\eta$ dynamically to ensure that (6) is satisfied. If (6) isn’t satisfied, we abort the step; i.e., we reduce $\eta$, and try again. If (6) is satisfied we accept the step. This is described in Algorithm 2. Note that in Algorithm 2 $\bar{\eta}\geq\frac{1}{\|A\|_{2}}$ holds always, and from this one can show the resulting step size $\eta\geq\frac{1-o(1)}{\|A\|_{2}}$ holds as $k\rightarrow\infty$.
Adaptive restarts.
In PDLP, we adaptively restart the PDHG algorithm in each outer iteration. The key quantity to our restarts at the $n$-th outer iteration is the normalized duality gap at $z$ which for any radius $r\in(0,\infty)$ is defined by
$\rho_{r}^{n}(z):=\frac{1}{r}\max_{(\hat{x},\hat{y})\in\{\hat{z}\in Z:\|\hat{z}-z\|_{\omega^{n}}\leq r\}}\{\mathcal{L}(x,\hat{y})-\mathcal{L}(\hat{x},y)\},$
introduced by [8]. Unlike the standard duality gap, $\max_{(\hat{x},\hat{y})\in Z}\{\mathcal{L}(x,\hat{y})-\mathcal{L}(\hat{x},y)\},$ the normalized duality gap is always a finite quantity. Furthermore, for any value of $r$ and $\omega^{n}$, the normalized duality gap $\rho_{r}^{n}(z)$ is 00 if and only if the solution $z$ is an optimal solution to (3) [8]; thus, it provides a metric for measuring progress towards the optimal solution. The normalized duality gap is computable in linear time [8]. For brevity, define $\mu_{n}(z,z_{\text{ref}})$ as the normalized duality gap at $z$ with radius $\|z-z_{\text{ref}}\|_{\omega^{n}}$, i.e.,
$\mu_{n}(z,z_{\text{ref}}):=\rho_{\|z-z_{\text{ref}}\|_{\omega^{n}}}^{n}(z),$
where $z_{\text{ref}}$ is a user-chosen reference point.
Choosing the restart candidate. To choose the restart candidate $z_{\text{c}}^{n,t+1}$ we call
[…]
This choice is justified in Remark 5 of [8].
Restart criteria. We define three parameters: $\beta_{\text{necessary}}\in(0,1)$, $\beta_{\text{sufficient}}\in(0,\beta_{\text{necessary}})$ and $\beta_{\text{artificial}}\in(0,1)$. In PDLP we use $\beta_{\text{necessary}}=0.9$, $\beta_{\text{sufficient}}=0.1$, and $\beta_{\text{artificial}}=0.5$. The algorithm restarts if one of three conditions holds:
  • (i)
    (Sufficient decay in normalized duality gap) $\mu_{n}(z_{\text{c}}^{n,t+1},z^{n,0})\leq\beta_{\text{sufficient}}\mu_{n}(z^{n,0},z^{n-1,0})\ ,$
  • (ii)
    (Necessary decay + no local progress in normalized duality gap)
    $\mu_{n}(z_{\text{c}}^{n,t+1},z^{n,0})\leq\beta_{\text{necessary}}\mu_{n}(z^{n,0},z^{n-1,0})\quad\text{and}\quad\mu_{n}(z_{\text{c}}^{n,t+1},z^{n,0})>\mu_{n}(z^{n,t}_{c},z^{n,0})\ ,$
  • (iii)
    (Long inner loop) $t\geq\beta_{\text{artificial}}k\ .$
The motivation for (i) is presented in [8]; it guarantees the linear convergence of restarted PDHG on LP problems. The second condition in (ii) is inspired by adaptive restart schemes for accelerated gradient descent where restarts are triggered if the function value increases [71]. The first inequality in (ii) provides a safeguard for the second one, preventing the algorithm from restarting every inner iteration or never restarting. The motivation for (iii) relates to the primal weights (Section 3). In particular, primal weight updates only occur after a restart, and condition (iii) ensures that the primal weight will be updated infinitely often. This prevents a bad choice of primal weight in earlier iterations from causing progress to stall for a long time.
Primal weight updates.
Define the map Combine$(\ell_{c},u_{c})$ which outputs a vector in $\mathbb{R}^{m}$ whose $i$th coordinate is given by $\max\left\{0,(\ell_{c})_{i}^{\star},(u_{c})_{i}^{\star}\right\}$ where for any $x\in\mathbb{R}\cup\{\pm\infty\}$, the operation $(x)^{\star}$ is equal to $|x|$ if it is finite or equal to zero otherwise. The primal weight is initialized using
[…]
where $\epsilon_{\text{zero}}$ is a small nonzero tolerance.
1 Function PrimalWeightUpdate($z^{n,0},z^{n-1,0},\omega^{n-1}$):
2       $\Delta_{x}^{n}=\|x^{n,0}-x^{n-1,0}\|_{2},\quad\Delta_{y}^{n}=\|y^{n,0}-y^{n-1,0}\|_{2}$ ;
3       if $\Delta_{x}^{n}>\epsilon_{\text{zero}}$ and $\Delta_{y}^{n}>\epsilon_{\text{zero}}$ then
4             return $\exp\left(\theta\log\left(\frac{\Delta_{y}^{n}}{\Delta_{x}^{n}}\right)+(1-\theta)\log(\omega^{n-1})\right)$
5       else
6             return $\omega^{n-1}$ ;
7
8       end if
9
10
Algorithm 3 Primal weight update
Algorithm 3 aims to choose the primal weight $\omega^{n}$ such that distance to optimality in the primal and dual is the same, i.e., $\|(x^{n,t}-x^{\star},\mathbf{0})\|_{\omega^{n}}\approx\|(\mathbf{0},y^{n,t}-y^{\star})\|_{\omega^{n}}$. By definition of $\|\cdot\|_{\omega}$,
$\|(x^{n,t}-x^{\star},\mathbf{0})\|_{\omega^{n}}=\omega^{n}\|x^{n,t}-x^{\star}\|_{2},\quad\|(\mathbf{0},y^{n,t}-y^{\star})\|_{\omega^{n}}=\frac{1}{\omega^{n}}\|y^{n,t}-y^{\star}\|_{2}.$
Setting these two terms equal yields $\omega^{n}=\frac{\|y^{n,t}-y^{\star}\|_{2}}{\|x^{n,t}-x^{\star}\|_{2}}$. Evidently, the quantity $\frac{\|y^{n,t}-y^{\star}\|_{2}}{\|x^{n,t}-x^{\star}\|_{2}}$ is unknown beforehand, but we attempt to estimate it using $\Delta^{n}_{y}/\Delta^{n}_{x}$. However, the quantity $\Delta^{n}_{y}/\Delta^{n}_{x}$ can change wildly from one restart to another, causing $\omega^{n}$ to oscillate. To dampen variations in $\omega^{n}$, we first move to a log-scale where the primal weight is symmetric, i.e., $\log(1/\omega^{n})=-\log(\omega^{n})$, and perform an exponential smoothing with parameter $\theta\in[0,1]$. In PDLP, we use $\theta=0.5$.
There are several important differences between our primal weight heuristic and proposals from Goldstein et al. [40, 41]. For example, Goldstein et al. make relatively small changes to the primal weights at each iteration, attempting to balance the primal and dual residual. These changes have to be diminishingly small because, in our experience, PDHG may be unstable if they are too big. In contrast, in our method, the primal weight is only updated during restarts, which in practice allows for much larger changes without instability issues. Moreover, our scheme tries to balance the weighted distance traveled in the primal and dual rather than the residuals.
Diagonal Preconditioning.
Preconditioning is a popular heuristic in optimization for improving the convergence of first-order methods. To avoid factorizations, we only consider diagonal preconditioners. Our goal is to rescale the constraint matrix $A$ to $\tilde{A}=D_{1}AD_{2}$ with positive diagonal matrices $D_{1}$ and $D_{2}$, so that the resulting matrix $\tilde{A}$ is “well balanced”. Such preconditioning creates a new LP instance that replaces $A,c,(\ell_{c},u_{c}),$ and $(\ell_{v},u_{v})$ in (1) with $\tilde{A}$, $\tilde{x}=D_{2}^{-1}x$, $\tilde{c}=D_{2}c$, $(\tilde{\ell_{c}},\tilde{u_{c}})=D_{1}(\ell_{c},u_{c})$, and $(\tilde{\ell_{v}},\tilde{u_{v}})=D_{2}^{-1}(\ell_{v},u_{v}).$ Common choices for $D_{1}$ and $D_{2}$ include:
  • Pock-Chambolle [75]: Pock and Chambolle proposed a family of diagonal preconditioners [^3] Diagonal preconditioning is equivalent to changing to a weighted $\ell_{2}$ norm in the proximal step of PDHG (weight defined by $D_{2}$ and $D_{1}$ for the primal and dual respectively). Pock and Chambolle use this weighted norm perspective. for PDHG parameterized by $\alpha$, where the diagonal matrices are defined by $(D_{1})_{jj}=\sqrt{\|A_{j,\cdot}\|_{2-\alpha}}$ for $j=1,...,m_{1}+m_{2}$ and $(D_{2})_{ii}=\sqrt{\|A_{\cdot,i}\|_{\alpha}}$ for $i=1,...,n$. We use $\alpha=1$ in PDLP. This is the baseline diagonal preconditioner in the PDHG literature.
  • Ruiz [80]: In an iteration of Ruiz scaling, the diagonal matrices are defined as $(D_{1})_{jj}=\sqrt{\|A_{j,\cdot}\|_{\infty}}$ for $j=1,...,m_{1}+m_{2}$ and $(D_{2})_{ii}=\sqrt{\|A_{\cdot,i}\|_{\infty}}$ for $i=1,...,n$. Ruiz [80] shows that if this rescaling is applied iteratively, the infinity norm of each row and each column converge to $1$.
For the default PDLP settings, we apply a combination of Ruiz rescaling [80] and the preconditioning technique proposed by Pock and Chambolle [75]. In particular, we apply $10$ iterations of Ruiz scaling and then apply the Pock-Chambolle scaling.
Infeasibility detection.
Inspired by [9], PDLP periodically checks whether any of three sequences—the difference of consecutive iterates, the normalized iterates, or the normalized running average—can serve as approximate certificates of infeasibility. We evaluate running averages starting from the most recent restart. Computing these sequences imposes little additional overhead, allowing PDLP to detect infeasibility efficiently. Notably, verifying all three sequences is often faster than relying on just one, since their convergence speed varies with the problem’s geometry [9].
Theoretical guarantees for the above enhancements.
While PDLP’s enhancements are motivated by theory, some of them may not preserve theoretical guarantees: we do not have a proof of convergence for the adaptive step size rule or primal weight update scheme. However, Applegate et al. [8] showed in a simpler setting with fixed primal weight and step size that adaptive restarts preserve convergence guarantees. Similarly, the theoretical gurantees for infeasibility detection [9] do not account for restarts.
4 Quickly finding approximately optimal but feasible solutions
In integer programming, it is typical for approximately optimal but feasible solutions to be acceptable. For example, branch and bound methods are commonly stopped as soon as the optimality gap falls below a predefined threshold (e.g., 1%) to ensure the user does not have to wait too long to receive a high-quality solution [69]. This pragmatic termination criteria motivates a new heuristic that we have developed that allows first-order methods to similarly find solutions with extremely small violation feasibility violations and moderate (e.g., 1%) duality gaps.
There are two key insights behind our heuristic:
  1. 1.
    Restarted PDHG tends to converge faster for feasibility problems than optimality problems. By a feasibility problem, we mean a problem with no objective, i.e.,
[…]
    subject to: $~{}\ell_{c}\leq Ax\leq u_{c}$
    $~{}\ell_{v}\leq x\leq u_{v},$
    Empirically, these problems can be orders of magnitude faster for PDHG to solve compared the original linear program (with nonzero objective).
  2. 2.
    Given a starting solution, restarted PDHG will converge to a nearby optimal solution. PDHG has the property that the distance (in terms of the PDHG norm) from the iterates to any optimal solution is nonincreasing [46]. Thus, if we warm start (7) from a primal solution obtained by applying PDHG to the original problem, we have reason to expect that the objective value will not substantially degrade, despite the fact that (7) ignores the objective function. Hence, if we have an approximately optimal solution obtained by PDHG, the result of solving (7) may retain the high quality objective and also be primal feasible to tight numerical tolerances. We can repeat this heuristic argument on the dual feasibility problem
    $\mathop{\rm maximize}_{y\in{\mathbb{R}}^{m},r\in{\mathbb{R}}^{n}}$ 00 0 (8)
    subject to: $c-A^{\top}y=r$
    $y\in\mathcal{Y}$
    $r\in\mathcal{R}.$
    to obtain a primal and dual solution that are almost exactly feasible with a small, but relatively larger, optimality gap.
These two insights are employed to design Algorithm 4. In particular, Algorithm 4 strives to find a feasible solution (i.e., $10^{-8}$ maximum constraint violation for both the primal and dual) with a relative duality gap less than the user-defined tolerance $\epsilon_{\text{rel-gap}}>0$, i.e.,
|$\frac{\absolutevalue{c^{\top}x+p(y;\ell_{c},u_{c})+p(r;\ell_{v},u_{v})}}{\max\{\absolutevalue{c^{\top}x},\absolutevalue{p(y;\ell_{c},u_{c})+p(r;\ell_{v},u_{v})}\}}\leq\epsilon_{\text{rel-gap}}.$
Appendix B gives a theoretical basis for these ideas. In particular, we show that PDHG with a fixed frequency restart scheme can achieve an $O(\beta/H(A)\log(1/\epsilon))$ iteration bound on LP instances where $\beta$ is its largest singular value of the constraint matrix $A$ and $H(A)$ is its Hoffman constant [48]. In other words, the new theory shows a runtime bound that is a constant independent of the right hand side or objective vector. In contrast, [8] shows an $O(\beta/H(K)\log(1/\epsilon))$ iteration bound for restarted PDHG on general linear programs where $K$ is a matrix containing the $A$ matrix, objective vector and right hand side. Moreover, some dependence on the right hand side and/or objective vector is fundamental to restarted PDHG: indeed there are examples where $H(A)=1$ but restarted PDHG is slow to converge (see, [47, Table 1]).
We note that similar arguments could be applied to other first-order methods, for example, replacing PDHG with AGD and minimizing $\|Ax-b\|^{2}$ s.t. $x\geq 0$ or using an adaptive restart scheme (following the techniques of [8]). As our main focus is computational, for simplicity we analyze the setting of PDHG with fixed frequency restarts.
  1. 1.
    Run PDLP on the LP instance. After completing $k$ iterations, for $k=100,200,400,800,\dots$, if the relative duality gap is less than $\epsilon_{\text{rel-gap}}$, pause “normal” PDLP iterations. Let $(x^{k},y^{k})$ be the average iterate (since the last restart) after that $k$-th iteration.
  2. 2.
    Run PDLP on the primal feasibility problem, i.e., Equation 7, starting from $(x^{k},0)$ with the initial primal weight and step size equal to the current values from Step 1. Run for $k/8$ iterations or until the primal constraint violation is less than the termination tolerance. In the former case return to Step 1, continuing “normal” PDLP iterations. Otherwise, let $\tilde{x}$ be output primal solution.
  3. 3.
    Run PDLP on the dual feasibility problem, i.e., Equation 8, starting from $(0,y^{k})$ with the initial primal weight and step size equal to the current values from Step 1. Run for $k/8$ iterations or until the dual constraint violation is less than the termination tolerance. Let $\tilde{y}$ be the output dual solution.
  4. 4.
    Check if $(\tilde{x},\tilde{y})$ meets the termination criteria, if not return to Step 1, resuming “normal” PDLP iterations.
Note that operations that are triggered periodically based on iteration counts (such as termination checks, adaptive restarts — see the paragraph “Periodic checks of restarts and termination criteria” in Section 3) take place in Steps 1, 2, and 3 using the iteration count assoiated with the current Step.
Algorithm 4 Feasibility polishing algorithm.
5 Implementation
The PDLP implementation we refer to in this work is a C++ implementation that is distributed as part of Google’s open-source OR-Tools library. This implementation of PDLP was first released with OR-Tools version 9.3 in March 2022. In this section, we briefly discuss the motivations and notable differences between this code and the Julia-based prototype used in the conference version of this work.
The primary motivation for a C++ implementation was the desire to integrate PDLP into production applications at Google, where concerns about technical uniformity make C++ more practical to deploy than Julia. Secondarily, the new implementation provided an opportunity to improve efficiency and generality, given that the Julia implementation was intended as a research prototype.
In addition to its improved accessibility thanks to its integration within both the OR-Tools framework and external modeling systems such as CVXPY [31], the C++ version has the following notable differences over the Julia code:
Feasibility polishing.
The Julia code did not have a feasibility polishing routine.
Multithreading.
The implementation parallelizes the computation of projections, norms, distances, and matrix-vector multiplications by the constraint matrix $x\mapsto Ax$ and its transpose $y\mapsto A^{\top}y.$ Any fixed number of threads is supported. In order to parallelize these computations, it divides the data (matrices or vectors) into shards and distributes the shards among threads. Each shard is then handled independently and then combined. By default, the number of shards is taken to be four times the number of threads.
More general LP form.
As mentioned in Section 2, we consider a more general LP format, commonly adopted by LP solvers, where inequalities are permitted to have both lower and upper bounds. Equalities are specified by setting the lower bounds equal to the upper bounds. The C++ code supports this formulation, in contrast to the Julia code that supports only single-sided inequalities and equalities.
Presolve.
The C++ implementation offers integrated presolving, in contrast with the conference version [7] where, for research purposes, we manually applied an external presolver. We did not develop new presolving routines; rather, we reused them from GLOP, the simplex-based LP solver included in OR-Tools. As a result of reusing GLOP’s presolving, the presolving code is single-threaded and not necessarily suitable for the largest scale of instances that PDLP is otherwise capable of handling. Thus, due to the size of the instances in the numerical experiments, we do not enable presolve in this work.
6 Large-scale benchmark instances
As part of this work, we developed and released a new test suite of large-scale LP instances. The set of problems is designed to be challenging for, if not beyond the capabilities of, today’s state-of-the-art solvers, yet small enough to fit in memory of top-end machines. The smallest problem in our new test suite has more than 125 million nonzeros, while the largest has over 6.3 billion. The collection is intended to be both realistic and diverse, but we also acknowledge that this collection is not comprehensive. The eleven LP instances are detailed in Table 1 and Table 2. Many of these problem instances were generated by the authors for the purposes of this project. There are few publicly available test instances of such a large scale, and it is our hope that these instances will be useful for other researchers.
The problems were scaled such that
  1. (i)
    The objective vector has values $1$, 00 or $-1$.
  2. (ii)
    The right-hand side has values $1$, 00 or $-1$, except in the case that a constraint has two nonzero finite, nonmatching lower and upper bounds. In this case we rescale such that the maximum absolute value of the right hand side is one.
This rescaling is implemented in the rescale_instance function in utils.jl.
All problems (after scaling) are available for download at https://www.oliverhinder.com/large-scale-lp-problems, and some have generators available along with detailed descriptions of the problems in the repository https://github.com/ohinder/large-scale-LP-test-problems as indicated in the right column of Table 1.
Instance name Brief description Full details available?
design-match Statistical matching with covariate balancing constraints for making causal inference on observational data. Loosely inspired by [92]. Yes
tsp-gaia-10m An LP lower bound for a travelling salesman problem. The underlying TSP instance [29] is the 10 million stars nearest to Earth in the Gaia Data Release 2 star catalog [21]. The variables correspond to 60,601,996 edges heuristically chosen to minimize the reduced-cost penalty when pricing the edges from the complete graph. The constraints are the 10,000,000 degree constraints and 7,016,681 cutting planes collected from extensive runs of the Concorde TSP solver [6] on sub-instances. No
tsp-gaia-100m Same as tsp-gaia-10m, except using the 100 million stars nearest to earth [28], with 1,184,557,727 edges, 100,000,000 degree constraints, and 62,934,799 cutting planes. No
heat-source-easy This problem considers a material with several randomly located heat sources. Given temperature measurements in at various parts of the material and a list of possible locations of the heat sources the aim is to recover the temperature distribution of the material. This is a PDE constrained optimization problem with linear partial differential equation constraints. Yes
heat-source-hard Same as heat-source-easy but significantly fewer measurement locations, more heat sources and more possible heat source locations. This makes the problem harder from a recover standpoint. Moreover, we find this problem is much harder for first-order methods than heat-source-easy. Yes
production-inventory A robust production-inventory problem [15]. Yes
qap-tho-150 An LP relaxation [4] of the tho150 quadratic assignment problem from QAPLIB [3]. Yes
qap-wil-100 An LP relaxation [4] of the wil100 quadratic assignment problem from QAPLIB [3]. Yes
world-shipping An LP relaxation of a MIP formulation of the LINERLIB [20] World instance. No
mediterranean-shipping An LP relaxation of a MIP formulation of the LINERLIB [20] Mediterranean instance. No
supply-chain A multicommodity flow problem motivated by optimizing the supply chain of a large retailer. Yes
Table 1: Descriptions of the LP problem instances.
problem name variables constraints nonzeros
design-match 40,000,000 22,000,135 2,760,000,000
tsp-gaia-10m 60,601,996 17,016,681 475,701,996
tsp-gaia-100m 1,184,557,727 162,934,799 6,337,834,450
heat-source-easy 31,628,008 15,625,000 125,000,000
heat-source-hard 31,628,008 15,625,000 125,000,000
production-inventory 18,270,600 4,650,850 500,049,700
qap-tho-150 249,783,750 6,705,300 1,005,795,000
qap-wil-100 49,015,000 1,980,200 198,020,000
world-shipping 228,867,510 15,304,282 688,658,522
mediterranean-shipping 208,479,461 7,490,593 628,927,462
supply-chain 201,000,100 2,210,100 403,000,100
Table 2: Problem instance sizes.
7 Numerical experiments
In this section we present numerical results on a new test suite of large-scale LP instances that we have developed and released as part of this work. Our goal is to show that first-order solvers, in particular PDLP with feasibility polishing, can provide feasible solutions to tight tolerances that are close-to-optimal, in reasonable time and without requiring enormous amounts of memory. We will compare PDHG with and without feasibility polishing to three commercial state-of-the-art Gurobi solvers: a barrier (i.e., interior point) method, and primal and dual simplex methods.
Note: we do not compare against other first-order methods or GPU implementations. We do not compare against first-order methods, because in the conference version of this paper we tested against SCS, a state-of-the-art first-order method and obtained almost an order of magnitude speed-up and do not anticipate it being competitive in our tests. We do not compare against GPU implementations because the release of the PDLP C++ code in 2022 pre-dated the GPU implementation of PDLP [56]. Additionally, some of the instances that we test exceed the RAM limit of a single GPU. We expect that a GPU implementation would outperform the OR-Tools code in wall clock time for instances that fit in the GPU’s memory.
Machines used
We used the Pitt computing cluster [2] to perform our experiments. There were two types of machines that we used: (i) an AMD EPYC 7302 (Rome) processor with 16 cores (32 threads) and 256 GB of memory and (ii) an Intel Xeon Platinum 8352Y (Ice Lake) processor with 32 cores (64 threads) and 1 TB of memory. The high memory machine used two and a half times more computing credits per hour so to limit its use we first ran a solver on an instance on the machine with 256 GB of memory and then only switched to the 1 TB machine if insufficient memory was avaliable. We estimate that our experiments would have cost around \$20,000 to run on Google cloud (excluding the cost of building and rescaling the instances). We ran PDLP with 16 threads on the 256 GB machine and 32 threads on the 1 TB machine. For Gurobi, we left the thread parameter at its default value, which means that it will automatically use as many threads as the machine has up to the number of virtual processors avaliable in the machine [1].
Multithreading performance of PDLP
From Figure 1 we can see that PDLP enjoys significant speed-up from more threads on large instances. With $32$ threads PDLP has an almost 8×8$\times$8 × speed-up on the design-match instance, one of the largest in the set. This speed-up comes from parallelizing the sparse matrix-vector operation and vector-vector operations. For smaller instances, the overall speed-up can be much less, in fact, more threads can even be counterproductive [56, Table 8]. We have identified two areas to improve our multithreading performance: thread synchronization and inefficient use of memory bandwidth. Implementing ideas from the literature on multithreaded sparse matrix-vector (SpMV) operations (e.g., [22, 51]) would lead to further improvements.
Figure 1: Speed up versus number of threads for PDLP on the design-match linear program on an AMD EPYC 7302 (Rome) processor with 16 cores (32 threads). We recorded the wall clock time to run 10,000 iterations with feasibility polishing disabled (1 measurement per number of threads).
Gurobi
We use Gurobi version 11.0.2 with crossover disabled.
Termination criteria
PDLP terminates when the maximum $\ell_{\infty}$ norms of the primal and dual residuals are less than $10^{-8}$, and the relative duality gap is less than $1\%$, i.e.,
|$\frac{\absolutevalue{c^{\top}x+p(y;\ell_{c},u_{c})+p(r;\ell_{v},u_{v})}}{\max\{\absolutevalue{c^{\top}x},\absolutevalue{p(y;\ell_{c},u_{c})+p(r;\ell_{v},u_{v})}\}}\leq 10^{-2}.$
We used the default Gurobi termination criteria, which after their own custom scaling, has an $\ell_{\infty}$ tolerance on the primal and dual residuals of $10^{-6}$, and objective gap tolerance of $10^{-8}$. Note that we do not believe that matching Gurobi’s termination criteria to ours would make a significant difference to the results. In particular, Gurobi’s barrier implementation mostly experiences OOM errors, and so is unaffected by the termination criteria; and simplex methods only find a primal and dual feasible solution at the optimum.
Results
Table 3 shows the time in hours taken per instance for each method to solve the problems and whether the method required a high memory machine or ran out of memory (shading the cell red). From this table one can see that Gurobi barrier only solves three instances because it hits memory limits in all but three instances, whereas PDLP never exceeds 1 TB of memory and only in one instance requires an excess of 256 GB. Gurobi’s primal and dual simplex is competitive with PDLP in terms of memory requirements but slow: combined dual simplex only solves four instances within the time limit of 144 hours. Finally, we see that feasibility polishing is extremely effective: PDLP without polishing only solves two instances, PDLP with polishing solves eight of eleven instances. Table 4 shows that the relative duality gap achieved by PDLP with feasibility polishing is typically much better than the $1\%$ threshold that we targeted. In particular, in all cases it obtains duality gap of less than $0.4\%$ and other than the two shipping instances the duality gap is always less than $0.1\%$.
Table 3 shows both PDLP and Gurobi had errors. Both of Gurobi’s errors gave the error code 10025. We believe this is a bug in Gurobi as the Gurobi error codes [42] state that the ‘user has called a query routine on a model with more than 2 billion non-zero entries’ even though we did not call a query routine. We have made Gurobi aware of this issue. Both of PDLP’s errors were numerical errors that occurred within a few hundred iterations.
PDLP Gurobi
Instance name Without polishing With polishing Barrier Dual Primal
design-match 68.0 9.3 OOM 75.3 32.4
tsp-gaia-10m TIME 3.0 28.1 90.1 40.2
tsp-gaia-100m TIME 21.1 OOM OOM OOM
heat-source-easy TIME 60.0 OOM ERR TIME
heat-source-hard TIME TIME OOM ERR TIME
production-inventory TIME TIME 5.8 TIME 43.7
qap-tho-150 ERR ERR OOM TIME TIME
qap-wil-100 44.4 0.3 OOM TIME TIME
world-shipping TIME 53.8 OOM TIME TIME
mediterranean-shipping TIME 32.6 OOM TIME TIME
supply-chain TIME 18.9 2.5 4.2 TIME
Table 3: Comparision of solvers in terms of the number of hours to solve instances (excluding time to read the instance). Termination criteria for Gurobi was set as default; termination criteria for PDLP was $10^{-8}$ maximum violation on primal and dual feasibility but $10^{-2}$ on the relative duality gap. Cells shaded red indicate runs that required the 1 TB RAM machine. TIME indicates that the solver exceeded 144 hours. OOM stands for ‘out of memory’, meaning the solver exceeded the 1 TB RAM limit. ERR denotes that there was an error during the solve.
Instance name Primal objective Absolute gap Relative gap (%)
design-match $2.87\times 10^{6}$ $5.53\times 10^{-2}$ $0.0000001$
tsp-gaia-10m $1.3\times 10^{8}$ $3.01\times 10^{3}$ $0.0012$
tsp-gaia-100m $1.94\times 10^{9}$ $5.82\times 10^{4}$ $0.0015$
heat-source-easy $5.16\times 10^{7}$ $9.06\times 10^{-4}$ $0.00000000088$
qap-wil-100 $2.21\times 10^{5}$ $1.14\times 10^{2}$ $0.026$
world-shipping $-1.18\times 10^{5}$ $3.7\times 10^{2}$ $0.16$
mediterranean-shipping $-1.07\times 10^{3}$ $6.95\times 10^{0}$ $0.33$
supply-chain $1.08\times 10^{8}$ $6.67\times 10^{3}$ $0.0031$
Table 4: Objectives and duality gaps produced from PDLP with feasibility polishing for instances that successfully terminated.
Acknowledgements
We thank Ed Klotz for his helpful discussion regarding the Irish Electricity instance and Joey Huchette for his feedback on an early draft of the paper.
This research was supported in part by the University of Pittsburgh Center for Research Computing, RRID:SCR_022735, through the resources provided. Specifically, this work used the SMP cluster, which is supported by NSF award number OAC-2117681.
Author contributions
Authors are ordered alphabetically. All authors contributed to writing, and regular discussion in meetings. The listed contributions span both the conference and journal papers.
DA and WS jointly designed the feasibility polishing implementation, and implemented most of the C++ code (except for the parts specifically mentioned elsewhere).
DA led the project after ML left Google, and generated the large-scale TSP instances which were our first large-scale test problems.
MD analyzed and implemented infeasibility detection features in both Julia and C++. He implemented and evaluated our step size choice against Malitsky and Pock’s. He also proofread and evaluated the large instances designed by OH.
OH devised the restart and primal weight scheme, implementing them in FirstOrderLP.jl; he proposed, and demonstrated the feasibility polishing concept. He also created the design-match, supply-chain and heat-source test problems. DA and WS ran preliminary experiments for this paper and OH ran the final experiments.
HL’s internship provided the impetus and initial code to pursue the idea of using first-order methods for LP. After his internship, HL, jointly with BOD, proposed the use of PDHG as the base algorithm. He designed and implemented the preconditioning scheme in both the FirstOrderLP.jl and C++ code. He also contributed to the development of the theory for restart strategies and infeasibility detection which influenced the implementation, and he implemented the production-inventory problem.
ML led the project while he was at Google, recruited and supervised HL, OH, and MD for their work on the project that was conducted at Google, contributed significantly to the FirstOrderLP.jl and OR-Tools implementations, created a framework for running large-scale experiments, suggested many aspects of the design of the numerical experiments, pushed forward the presolve enhancement, and led the open sourcing effort.
BOD, jointly with HL, proposed the usage of PDHG as the core algorithm to apply to LP. He also provided advice and feedback on the restarting techniques and the usage of SCS, and helped with data analysis and figure generation for the NeurIPS paper.
WS performed the initial experiments demonstrating the effectiveness of mirror-prox and PDHG for LP including some initial heuristics that inspired the primal weight scheme and restarts, and designed and implemented the sharded parallelism mechanism and the adaptive step size scheme.
Appendix A Reduced cost selection for PDLP
As discussed in Section 2, there are multiple choices for recovering the reduced costs when running PDLP. These choices impact convergence detection but not the sequence of iterates. A natural choice is to set
$r\leftarrow\mathop{\bf proj}_{\mathcal{R}}(c-A^{\top}y);$
however, an alternative that we found to be useful is to set
$r_{i}\leftarrow\begin{cases}c_{i}-a_{i}^{\top}y&\text{if }c_{i}-a_{i}^{\top}y>0\text{ and }x_{i}-(\ell_{v})_{i}\leq\absolutevalue{x_{i}}\\ c_{i}-a_{i}^{\top}y&\text{if }c_{i}-a_{i}^{\top}y<0\text{ and }(u_{v})_{i}-x_{i}\leq\absolutevalue{x_{i}}\\ 0&\text{otherwise}\end{cases}$
where $a_{i}$ is the $i$th column of $A$. This choice is more robust to artifically large bounds. For example, a relatively small residual reduced cost when a variable bound is extremely large, e.g., $10^{12}$ could prevent convergence through its effect on the dual objective value. This more advanced choice is only used when we run PDLP without feasibility polishing.
Appendix B Theoretical basis for feasibility polishing
For simplicity, we assume the LP instance takes the form $\min c^{\top}x$ such that $Ax=b,x\geq 0$ and focus on primal feasibility polishing (i.e., $c=\mathbf{0}$). However, this analysis generalizes to other settings such as the LP format described in Section 2 and dual feasibility polishing.
In this simple case, PDHG reduces to
$x^{t+1}$ $\leftarrow\mathop{\bf proj}_{x\geq 0}\left(x^{t}-\eta(c-A^{\top}y^{t})\right)$ (9a)
$y^{t+1}$ $\leftarrow y^{t}-\eta(b-A(2x^{t+1}-x^{t}))\ .$ (9b)
Let $X_{F}:=\{x\geq 0:Ax=b\}$ be set of feasible solutions. Let $\mathop{\bf dist{}}(X,x):=\inf_{z\in X}\|x-z\|_{2}$. The Hoffman constant of $A$ is the largest constant, $H(A)$, such that
$H(A)\mathop{\bf dist{}}(X_{F},x)\leq\|Ax-b\|_{2}\quad\quad\text{for all }b\text{ such that }X_{F}\neq\emptyset.$ (10)
Thanks to [48], we know that this constant is always positive.
Lemma 1.
Suppose that $c=\mathbf{0}$ (so we have a primal feasibility problem), $0<\eta<1/\|A\|_{2}$ and $y^{0}=\mathbf{0}$. Then PDHG, i.e., (9) with $\bar{x}^{t}=\frac{1}{t}\sum_{i=1}^{t}x^{i}$ satisfies
$\|\bar{x}^{t}-x^{0}\|_{2}$ $\leq(q+2)\mathop{\bf dist{}}(x^{0},X_{F})$ (11)
$H(A)\mathop{\bf dist{}}(X_{F},\bar{x}^{t})$ $\leq\frac{(q+2)\mathop{\bf dist{}}(x^{0},X_{F})}{\eta t}$ (12)
where $q=4\frac{1+\eta\|A\|_{2}}{1-\eta\|A\|_{2}}$.
Proof.
Let $Z^{\star}$ denote the set of minimax solutions to $\min_{x\geq 0}\max_{y}y^{\top}Ax-b^{\top}y$. From Corollary 2 of [8] we have
$\|\bar{z}^{t}-z^{0}\|_{2}\leq(q+2)\mathop{\bf dist{}}(z^{0},Z^{\star})=(q+2)\mathop{\bf dist{}}(x^{0},X_{F})$ (13)
which by $\|\bar{x}^{t}-x^{0}\|_{2}\leq\|\bar{z}^{t}-z^{0}\|_{2}$ gives (11).
Let $\bar{z}^{t}=\frac{1}{t}\sum_{i=1}^{t}z^{i}$. By Corollary 2 of [8] with $C=1/\eta$ and the fact that $y^{0}=\mathbf{0}$ we have
$\rho_{\|\bar{z}^{t}-z^{0}\|_{2}}(\bar{z}^{t})\leq\frac{2\|\bar{z}^{t}-z^{0}\|_{2}}{t\eta}.$ (14)
Moreover, by [8] equation (22), we have
$\rho_{\|\bar{z}^{t}-z^{0}\|_{2}}(\bar{z}^{t})\geq\|A\bar{x}^{t}-b\|_{2}+\|(-c+A^{\top}\bar{y}^{t})^{+}\|_{2}\geq\|A\bar{x}^{t}-b\|_{2}$ (15)
where $+$ is the positive part operator. Combining (10), (13), (14), and (15) with this inequality yields (12). ∎
1
2Input: A starting point $x^{0,0}$, a step-size $\eta\in(0,\infty)$, a length for the restart period $\ell\in{\mathbb{N}}$;
3 Initialize outer loop counter $n\leftarrow 0$;
4 Initialize duals $y^{0,0}\leftarrow\mathbf{0}$;
5
6for $n=0,1,2,\dots$ do
7       initialize the inner loop. inner loop counter $t\leftarrow 0$ ;
8       repeat
9             Take a step of PDHG, i.e., calculate $(x^{n,t+1},y^{n,t+1})$ from $(x^{n,t},y^{n,t})$ via (9) ;
10             $\bar{x}^{n,t+1}\leftarrow\frac{1}{t+1}\sum_{i=1}^{t+1}x^{n,i}$;
11             $t\leftarrow t+1$;
12
13      until $t\geq\ell$ ;
14      restart the outer loop. $x^{n+1,0}\leftarrow\bar{x}^{n,t+1}$, $y^{n+1,0}\leftarrow\mathbf{0}$, $\tau^{n}\leftarrow\ell$, $n\leftarrow n+1$;
15
16 end for
Algorithm 5 A first-order method for quickly finding a nearby feasible solution
Theorem 1.
Suppose that $c=\mathbf{0}$ (so we have a primal feasibility problem). Consider Algorithm 5 with $0<\eta<1/\|A\|_{2}$ and $\ell\geq t^{\star}\vcentcolon=\Bigg{\lceil}\frac{2(q+2)}{\eta H(A)}\Bigg{\rceil}$ where $q=4\frac{1+\eta\|A\|_{2}}{1-\eta\|A\|_{2}}$. Then, for all $n\geq 1$
$\mathop{\bf dist{}}(x^{n,0},X_{F})$ $\leq 2^{-n}\mathop{\bf dist{}}(x^{0,0},X_{F})$ (16a)
$\|x^{n,0}-x^{0,0}\|_{2}$ $\leq(q+2)\mathop{\bf dist{}}(x^{0,0},X_{F}).$ (16b)
Proof.
We obtain (16a) from (12) and the assumption that $\ell\geq t^{\star}$. Observe that
$\|x^{k,0}-x^{0,0}\|_{2}$ $\leq\|x^{k,0}-x^{k-1,0}\|_{2}+\|x^{k-1,0}-x^{0,0}\|_{2}\leq(q+2)\mathop{\bf dist{}}(x^{k,0},X_{F})+\|x^{k-1,0}-x^{0,0}\|_{2}$ (17)
$\leq(q+2)2^{-k}\mathop{\bf dist{}}(x^{0,0},X_{F})+\|x^{k-1,0}-x^{0,0}\|_{2}$
where the first inequality uses the triangle inequality, the second inequality uses (11), and the last inequality uses (16a). Using (17) will prove the following induction hypothesis:
$\|x^{n,0}-x^{0,0}\|_{2}\leq\left(\sum_{i=1}^{n}2^{-i}\right)(q+2)\mathop{\bf dist{}}(x^{0,0},X_{F}).$
In particular, (17) with $k=1$ shows the hypothesis for $n=1$ since $x_{k-1,0}=x^{0,0}$. If the hypothesis holds for $n=k-1$ then by (17) it holds for $n=k$. Thus we have proved the induction hypothesis. Using the formula for the sum of an infinite geometric series we get $\sum_{i=1}^{n}2^{-i}\leq\sum_{i=1}^{\infty}2^{-i}=1$, thus the induction hypothesis implies (16b). ∎